OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation x^2/9-y^2/16 = 1

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Compare it to the standard form: \(\large \frac{x^2}{a^2} - \frac{y^2}{b^2} = 1\) and identify a and b.

OpenStudy (anonymous):

a=3 b=4

OpenStudy (aum):

Correct. Vertex at (-a, 0) and (a, 0) Focus at (-c, 0) and (c, 0) where c^2 = a^2 + b^2

OpenStudy (anonymous):

oh wow thank you!!

OpenStudy (aum):

You are welcome.